# CFT Constraints on Parity-odd Interactions with Axions and Dilatons

### v1

@@ 5 Physical implications for chiral/gravitational backgrounds
$\mathcal{L}_{axion}\supset\frac{\phi}{\Lambda}R_{\mu\nu}\tilde{R}^{\mu\nu}$ (58)
where $R_{\mu\nu}$ is the Ricci tensor. This correlator mediates an interaction between the pseudoscalar and two gravitational waves at semiclassical level and is part of the effective action generated by integrating out conformal matter in quantum corrections. The interaction is the analogue of the the $(\phi/f)F\tilde{F}$ interaction of ordinary axion-physics, which is associated with the $\langle JJO\rangle$ $(\Delta_{3}=4)$ in (39). This process, referred to as conformal backreaction [40], is based on the hypothesis that the universe underwent a conformal phase before any ordinary mechanism based on spontaneous symmetry breaking took place.
@@
The $\langle JJO\rangle$, where $J$ is a spin-1 current to gauge fields like the hypercharge $Y$ and axion-like fields, can also be linked to gravity nonperturbatively. In this case, $O$ interpolates with a trace anomaly of a stress-energy tensor of odd parity. This point has come to the attention of several groups in the recent literature. The apperance of this anomaly with Standard Model fermions has been attributed to the implementation of regularization procedure in the perturbative analysis of such contributions with conflicting results [41, 42, 43, 44, 45, 46, 47, 48].
In a previous analysis [15], we showed that the Conformal Ward Identities (CWIs) permit a minimal form of this interaction. However, such approach is nonperturbative. In this context, the axion/dilaton field $(\chi)$ would interact according to the form described in (58).
@@
A final comment concerns the $\langle TJO\rangle$ that vanished identically as an off-shell correlator in a CFT, both in its even and odd components (see also Appendix B). This result can be interpreted as a constrain on the absence of an axion/dilaton to a spin-1 mixing in the presence of an external gravitational field. This can be viewed as a constrain similar to the axion to photon transition in the presence of an external magnetic field, which is the basic process for axion detection with helioscopes [57].
@@ 6 Conclusions
We have analyzed all possible 3-point functions constructed with the energy-momentum tensor, currents and at least one scalar operator in CFT. Most of the correlators are constrained to be zero, except some specific case which are protected from vanishing by their anomalies.
In particular, all the correlators built with at least two scalar operator vanish because of Lorentz invariance. There is no need to require invariance under the full conformal group, as it is not even possible to construct tensorial structures for them.
@@ Acknowledgements
This work is partially funded by the European Union, Next Generation EU, PNRR project "National Centre for HPC, Big Data and Quantum Computing", project code CN00000013; by INFN, inziativa specifica QG-sky and by the grant PRIN 2022BP52A MUR "The Holographic Universe for all Lambdas" Lecce-Naples.
@@
The most general solution of the CWIs for our correlators can be written in terms of integrals involving a product of three Bessel functions, namely 3K integrals. In this appendix, we will illustrate such integrals and their properties. For a detailed review on the the topic, see also [2, 3, 33].
@@ A.1 Definition and properties
From this analysis, it is simple to relate the form factors to the 3K integrals. Indeed, the dilatation WI of each from factor tells us that this needs to be written as a combination of integrals of the following type
@@ A.2 Appell functions
In the case of scalar primary operators, for example, of scaling dimensions $\Delta_{i}$, and momenta $p_{1},p_{2},p_{3}$, the solutions expressed by 3K integrals can be directly related to the four Appell functions [1] characterised by 4 pairs of indices $(a_{i},b_{j})$ $(i,j=1,2)$. These are the indices that in the change of variables $(p_{1}^{2},p_{2}^{2},p_{3}^{2})\to(p_{1}^{2},x,y)$, $x=p_{2}^{2}/p_{1}^{2},$ $y=p_{3}^{2}/p_{1}^{2}$ reduce the special conformal constraints to Appell hypergeometric equations deprived of $1/x$ or $1/y$ singularities [32, 58]. Setting
@@
where $(\alpha,i)=\Gamma(\alpha+i)/\Gamma(\alpha)$ is the Pochammer symbol. We will refer to $\alpha\ldots\gamma^{\prime}$ as to the first,$\ldots$, fourth parameters of $F_{4}$.
The 4 independent solutions of the Appell system of equations are then all of the form $x^{a}y^{b}F_{4}$, linearly combined in a Bose-symmetric form. Specifically, the solution for the parity-even correlator with three scalar operators takes the general form
@@ A.3 Zero momentum limit
When solving the secondary CWIs, it may be useful to perform a zero momentum limit. In this subsection, we review the behaviour of the 3K integrals in the limit $p_{3}\rightarrow 0$. In this limit, the momentum conservation gives
@@ A.4 Divergences and regularization
In general, the regularisation parameters $u$ and $v_{i}$ are arbitrary. However, in certain cases, there may be some constraints on them.
@@
where $\beta_{t}=\beta_{1}+\beta_{2}+\beta_{3}$. All tensorial massless 1-loop 3-point momentum-space integrals can also be expressed in terms of a number of 3K integrals when their tensorial structure is resolved by standard methods (for the exact expressions in this case see Appendix A.3 of [2]).
Appendix B A comment on the $\langle TJO\rangle$
Here we add few more comments on the implications of the vanishing of the $\langle TJO\rangle$-mediated off-shell amplitudes, derived from the CFT analysis. We consider an off-shell (massive) graviton that couples to the $\langle TJO\rangle$. The full decomposition of the background gravitational into its spin-2, spin-1 and spin-0 fluctuations is given by
$h_{\mu\nu}=h_{\mu\nu}^{(2)}+h_{\mu\nu}^{(1)}+h_{\mu\nu}^{(0)}$ (85)
with
$h_{\mu\nu}^{(2)}=h_{\mu\nu}^{\text{TT}}$ (86)
where $h_{\mu\nu}^{\text{TT}}$ is the transverse-traceless part. The spin-1 component is expressed as
$h_{\mu\nu}^{(1)}=\partial_{\mu}D_{\nu}+\partial_{\nu}D_{\mu}$ (87)
where $D_{\mu}$ is a divergenceless vector field $(\partial^{\mu}D_{\mu}=0)$. The spin-0 component is given by
$h_{\mu\nu}^{(0)}=\left(\partial_{\mu}\partial_{\nu}-\frac{1}{4}\eta_{\mu\nu}\Box\right)\sigma+\frac{1}{4}\eta_{\mu\nu}h$ (88)
where $\sigma$ is a scalar potential and $h=h^{\mu}_{\mu}$ is the trace of the perturbation. We will denote with an index (±)plus-or-minus($\pm$)( ± ) the parity of the component. For example $h_{\mu\nu}^{(2)(+)}$ is the spin-2 component, of positive parity, as well as $h_{\mu\nu}^{(0)(+)}$. $h_{\mu\nu}^{(1)}$ does not have definite parity. The decomposition allows to extract helicities ${\bf h}=\pm 2,\pm 1,0$ for the spin-2 part, ${\bf h}=\pm 1,0$ for the spin-1 and ${\bf h}=0$ for the spin-0.
If we couple a massive spin-1 $(B_{0},B_{i})$ to the vector current $J$, the two components $B_{0}$ and $B_{i}$ carry positive and negative parities $(B_{0}^{(+)},B_{i}^{(-)})$ respectively. It is clear that conformal symmetry constrains such off-shell interaction in all of their components both for parity conserving and parity violating transitions. While some are trivially vanishing by helicity conservation, others are effectively set to vanish.

### v2

@@ 6 Physical implications for chiral/gravitational backgrounds
$\mathcal{L}_{axion}\supset\frac{\phi}{\Lambda}\varepsilon^{\mu\nu\rho\sigma}R_{\beta\mu\nu}^{\alpha}R_{\alpha\rho\sigma}^{\beta}$ (67)
This correlator mediates an interaction between the pseudoscalar and two gravitational waves at semiclassical level and is part of the effective action generated by integrating out conformal matter in quantum corrections. The interaction is the analogue of the $(\phi/f)F\tilde{F}$ interaction of ordinary axion-physics, which is associated with the $\langle JJO\rangle$ $(\Delta_{3}=4)$ in (48). This process, referred to as conformal backreaction [40], is based on the hypothesis that the universe underwent a conformal phase before any ordinary mechanism based on spontaneous symmetry breaking took place.
@@
The $\langle JJO\rangle$, where $J$ is a spin-1 current to gauge fields like the hypercharge $Y$ and axion-like fields, can also be linked to gravity non-perturbatively. In this case, $O$ interpolates with a trace anomaly of a stress-energy tensor of odd parity. This point has come to the attention of several groups in the recent literature. The appearance of this anomaly with Standard Model fermions has been attributed to the implementation of regularization procedures in the perturbative analysis of such contributions with conflicting results [41, 42, 43, 44, 45, 46, 47, 48].
In a previous analysis [15], we showed that the CWIs permit a minimal form of this interaction. However, such approach is non-perturbative. In this context, the axion/dilaton field $(\chi)$ would interact according to the form described in (67).
@@
A final comment concerns the $\langle TJO\rangle$ that vanished identically as an off-shell correlator in a CFT, both in its even and odd components. This result can be interpreted as a constraint on the absence of an axion/dilaton to a spin-1 mixing in the presence of an external gravitational field. This can be viewed as a constraint similar to the axion to photon transition in the presence of an external magnetic field, which is the basic process for axion detection with helioscopes [57].
@@ 7 Conclusions
We have analyzed all possible 3-point functions constructed with the energy-momentum tensor, currents and at least one scalar operator in CFT. Most of the correlators are constrained to be zero, except some specific cases which are protected from vanishing by their anomalies.
In particular, all the correlators built with at least two scalar operators vanish because of Lorentz invariance. There is no need to require invariance under the full conformal group, as it is not even possible to construct tensorial structures for them.
@@ Acknowledgements
This work is partially funded by the European Union, Next Generation EU, PNRR project "National Centre for HPC, Big Data and Quantum Computing", project code CN00000013; by INFN, iniziativa specifica QG-sky and by the grant PRIN 2022BP52A MUR "The Holographic Universe for all Lambdas" Lecce-Naples.
@@
The most general solution of the CWIs for our correlators can be written in terms of integrals involving a product of three Bessel functions, namely 3K integrals. In this appendix, we will illustrate such integrals and their properties. For a detailed review on the topic, see also [2, 3, 33].
@@ A.1 Definition and properties
From this analysis, it is simple to relate the form factors to the 3K integrals. Indeed, the dilatation Ward identity of each from factor tells us that this needs to be written as a combination of integrals of the following type
@@ A.2 Appell functions
In the case of scalar primary operators, for example, of scaling dimensions $\Delta_{i}$, and momenta $p_{1},p_{2},p_{3}$, the solutions expressed by 3K integrals can be directly related to the four Appell functions [1] characterized by four pairs of indices $(a_{i},b_{j})$ $(i,j=1,2)$. These are the indices that in the change of variables $(p_{1}^{2},p_{2}^{2},p_{3}^{2})\to(p_{1}^{2},x,y)$, $x=p_{2}^{2}/p_{1}^{2},$ $y=p_{3}^{2}/p_{1}^{2}$ reduce the special conformal constraints to Appell hypergeometric equations deprived of $1/x$ or $1/y$ singularities [32, 58]. Setting
@@
where $(\alpha,i)=\Gamma(\alpha+i)/\Gamma(\alpha)$ is the Pochhammer symbol. We will refer to $\alpha\ldots\gamma^{\prime}$ as to the first,$\ldots$, fourth parameters of $F_{4}$.
The four independent solutions of the Appell system of equations are then all of the form $x^{a}y^{b}F_{4}$, linearly combined in a Bose-symmetric form. Specifically, the solution for the parity-even correlator with three scalar operators takes the general form
@@ A.3 Zero momentum limit
When solving the secondary CWIs, it may be useful to perform a zero momentum limit. In this subsection, we review the behavior of the 3K integrals in the limit $p_{3}\rightarrow 0$. In this limit, the momentum conservation gives
@@ A.4 Divergences and regularization
In general, the regularization parameters $u$ and $v_{i}$ are arbitrary. However, in certain cases, there may be some constraints on them.
@@
where $\beta_{t}=\beta_{1}+\beta_{2}+\beta_{3}$. All tensorial massless 1-loop 3-point integrals can also be expressed in terms of a number of 3K integrals when their tensorial structure is resolved by standard methods (for the exact expressions in this case see Appendix A.3 of [2]).